Solve the following equation using the substitution method.
x^4 –5x^2 –14=0
Complete the following steps to solve the above equation:
Let u=x^2, substitute the variables, and write the new equation.
Now factor the new equation, apply the zero-product principle, and solve for u.
Now substitute u=x^2 to solve for x.

Answer by jsmallt9(3758)

The directions you were given are very good. The only improvements I can make are:
Instead of:
"Let u=x^2, substitute the variables, and write the new equation."
I would add
"Let u=x^2. Then u^2 = (x^2)^2 = x^4, substitute u and u^2 for x^2 and x^4"
and then instead of:
"Now substitute u=x^2 to solve for x."
I would say:
"Replace the u with x^2 and solve for x"